hey i need help making a graph for this problem. thanks!!
It costs a business $20 to make one compact disk player and $10 to make one radio. Each week the company must make a combined total of at least 50 CD players and radios. At least as many CD players as radios must be manufactured. Determine how many CD players and radios should be made to minimize weekly costs.

Answer by ajs12405(1)

x=number of cds
y=number of radios
x+y>50
x>=y
20x+10Y=C (cost minimize)
20x+10y=750
y=-.5x+75
x=25 y=25
20(25) +10(25)=750
20(26) +10(25)=770
You should produce 25 CD players and 25 radios to minimize costs at $750 per week.